Fermi National Laboratory
 

Operation Summary for December 22, 2010 - January 3, 2011

 

The Integrated luminosity for the period from 12/22/10 to 1/3/11 was 53.9 pb-1 (inverse picobarns). NuMI reported receiving 8.81E18 protons from 12/2710 to 1/3/11.

Wednesday Dec. 22

The day shift began with Operations monitoring Tevatron (TEV) store 8383 121.96E30) and the Recycler stash (346.9E10), with the Antiproton Source (Pbar) stacking (11E10), with the Linac, Booster, and Main Injector (MI), operating normally, with the NuMI, MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8383 at 9:16 AM and prepared the D0 collision hall for access; D0 came out of access at 11:22 AM. Operators put the TEV through an orbit check and then into HEP shot setup. Operations established store 8385 at 1:25 PM with an average initial luminosity (AIL) of 368.71E30.

During the evening shift, Operations established beam to the new Meson FTBF T932 (Diamond) experiment at 5:33 PM. T-932 ran until 6:39 PM. A Linac RF station (LRF5) developed a bad switch tube grid at 10:05 PM. Linac personnel began replacing the switch tubes.

Thursday Dec. 23

The midnight shift began with Operations monitoring store 8385 (103.24E30) and the stash (231.9E10), with Pbar stacking (18.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Linac resumed sending beam at 6:50 AM. Operations established beam to T-932 at 7:30 AM.

During the day shift, Operations terminated store 8385 at 10:16 AM, put the TEV into shot setup, and established store 8386 at 11:55 AM with an AIL of 314.47E30.

During the evening shift, Operations disabled beam to T-932 at 6:08 PM.

Friday Dec. 24

The midnight shift began with Operations monitoring store 8386 (82.37E30) and the stash (300E10), with Pbar stacking (9.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8386 at 4:44 AM, put the TEV into shot setup, and established store 8387 at 6:05 AM with an AIL of 326.55E30.

During the day shift, Operations stacked and transferred antiprotons to the Recycler as appropriate. T-932, NuMI, and MiniBooNE took beam when it was available.

During the evening shift, Operations terminated store 8387 at 9:10 PM, put the TEV into shot setup, and established store 8388 at 11:15 PM with an AIL of 350.2E30. A MI Lambertson (LAM62) and kicker (MI-62) magnet power supply tripped off, which caused some delays in shot setup. Operations disabled beam to T-932 shortly after 6 PM.

Saturday Dec. 25

The midnight shift began with Operations monitoring store 8388 (235.18E30) and the stash (69.9E10), with Pbar stacking (20.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported RF problems in MI and Booster; BRF11 tripped off on a bias supply ground fault and would not reset; operators brought BRF17 online in its place. Operators tuned Booster after the RF problems.

During the day shift, Operations reported more Booster RF problems that resulted in more tuning. Operations reported that the TEV sector B1 cryo system (downstream side) started going unstable, eventually resulting in a quench at B2 at 9:24 AM. Operators investigated and discovered that vacuum problems caused the instability. Vacuum technicians accessed and added extra turbo pumps, but this did not resolve the issue. Their leak checking pointed toward a nitrogen to cryostat vacuum leak. Operators put in a special timeline to slow down the stacking rate – the Recycler was near its stash cap limit. Operators conducted Booster and MI tuning.

During the evening shift, Operators monitored the accelerator complex.

Sunday Dec. 26

The midnight shift began with the TEV warm up continuing, with Operations monitoring the stash (453E10), with Pbar stacking but not stashing, and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TEV remained down. Operators continued to tune MI and Booster. Operations established beam to T-932 at 4:43 AM. Operations reported some minor downtime due to Booster and MI RF trips.

During the day shift, Operations reported that the TEV vacuum woes continued.

CHL personnel started warming up the B1 cryo system at 8:30 AM. Operators made various adjustments and continued tuning. Operators prepared the TEV A-E sectors for a Mechanical Support access into the Transfer Hall to stage equipment for repairs later.

During the evening shift, Operations reported that the TEV B1 cryo system warm up continued. Pbar stacked antiprotons, but did not transfer any to the Recycler. Operations disabled beam to T-932 at 10 PM.

Monday Dec. 27

The midnight shift began with the TEV warm up continuing, with Operations monitoring the stash (~440E10), with Pbar stacking (~100E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-932 at 4:22 AM, but turned off the beam at 7 AM for a Transfer Hall access. Operations reported some downtime due to an overheating crate power supply in MI (crate $6A), which operators replaced.

During the day shift, the TEV B1 warm up continued. Operations reported that Pbar stacked antiprotons until 9:28 AM, when operators put the machine into standby mode for necessary repairs. RF personnel reported that they tied off BRF11 from the anode power supply, which caused about 10 minutes of downtime. Personnel from Booster, MI and Recycler conducted studies through out the shift. There were no transfers of antiprotons to the Recycler during this shift.

During the evening shift, Pbar remained in standby all shift. The TEV warm up continued. Recycler studies concluded at 8 PM.

Tuesday Dec. 28

The midnight shift began with the TEV warm up continuing, with Operations monitoring the stash (303.7E10), with Pbar stacking (167.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TEV warm up continued. Pbar remained in standby. Operations turned off the Booster beam at 7 AM and prepared Booster and MI for access.

During the day shift, the Booster, MI, and TEV were all in access. B1 warm up continued. Operations had MI ramping by 11:43 AM. Booster came out of access at 1:42 PM. Pbar resumed stacking and the Recycler stashed antiprotons. Operations put Pbar back into standby after the stash and the Recycler began cooling studies.

During the evening shift, Recycler cooling studies continued until 8:47 PM. The Recycler also conducted extraction and chromaticity studies.

Wednesday Dec. 29

The midnight shift began with the TEV warm up continuing, with Operations monitoring the stash (17.3E10), with Pbar stacking (42.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TEV work continued. Pbar remained in stacking standby mode.

During the day shift, the TEV work continued. Operations reported that the B1 cryo system warm up had completed early on the shift and Mechanical Support personnel started vacuum leak checking. Operators turned off the TEV RF stations. Recycler and Pbar studies took place this shift. The NuMI absorber LCW power supply temperature alarmed; NuMI personnel changed out the filters and the temperature came back down. Operations reported some minimal downtime due to a couple of Linac LRF5 power amplifier trips.

During the evening shift, the TEV vacuum work continued. Vacuum techs reported finding a leak on a bellows between the spool piece and the turnaround box. They will replace the spool piece. Pbar and Recycler studies took place. Pbar went back to stacking standby at 8 PM. Mechanical Support personnel moved the new spool piece into place; once this access was complete, Switchyard (SY) personnel took beam to the dump for studies.

Thursday Dec. 30

The midnight shift began with the TEV down due to vacuum work at B1, with Operations monitoring the Pbar stack (30.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TEV vacuum work continued. A kicker (KPS5S) tripped off on a Thyratron power fault and would not reset. Kicker personnel came in to investigate and got it running again. Unfortunately, poor beam positions around the MI-8 collimators started pulling the MI beam permit; operators and MI personnel investigated.

During the day shift, Operations put Pbar into standby around 8:40 AM while a kicker (KPS5S) was being repaired. SY personnel conducted studies from 8:40 AM to 12:30 PM. Alignment personnel reported that they had finished their work on the new spool piece around 12:30 PM. Operations reported short downtimes due to MIRF17 & 18, LRF2, and a couple of south anode power supply trips.

During the evening shift, the TEV work continued. Operators reported that the Booster East Anode Power Supply flow meter tripped off; operators reset the meter.

Friday Dec. 31

The midnight shift began with the TEV down due to vacuum work at B1, with Operations monitoring the Pbar stack (30E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TEV B1 spool work continued. Operators tuned MI. Operators prepared the TEV transfer Hall for an access.

During the day shift, the TEV B1 spool work continued. Operations reported short downtimes due to Linac and Booster RF problems.

During the evening shift, the TEV B1 spool work continued. Operations reported that the NuMI dehumidifiers tripped off. NuMI personnel came in and got the dehumidifiers back on. LRF2 tripped off several times. BRF19 tripped off multiple times.

Saturday Jan. 1

The midnight shift began with the TEV down due to vacuum work at B1, with Operations monitoring the Pbar stack (29.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TEV B1 repairs continued. NuMI personnel re-enabled their condensate tank pump.

During the day shift, the TEV B1 repairs continued. Pbar maintained a stack of ~20 mA .

During the evening shift, the TEV B1 repairs continued. Operations monitored and tuned the Proton Source.

Sunday Jan. 2

The midnight shift began with the TEV down due to vacuum work at B1, with Operations monitoring the Pbar stack, and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TEV B1 repairs continued. Operations discovered BRF19 50 volt power supply interlock problems in Booster and switched over to BRF11. The MI Longitudinal Damper tripped off four times during the shift; it reset. Recycler personnel conducted Pbar “Drag Rate” studies. LLRF5 tripped off; it reset.

During the day shift, the TEV B1 repairs continued. Pelletron personnel conducted beam studies. Operators monitored the Pbar stack, ~28mA.

During the evening shift, the TEV B1 repairs continued. Recycler personnel conducted studies.

Monday Jan. 3

The midnight shift began with the TEV down due to vacuum work at B1, with Operations monitoring the Pbar stack, and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TEV B1 repairs continued. Operators prepared the Pbar Rings, Transport, Pre-Target, and Pre-Vault for an access at 5 AM. Operations switched off MI at 6 AM and made the area ready for an access by 7 AM. Operators had F-Sector ready for an access at 7:10 AM. Mechanical Support personnel localized the B13-5 single-phase leak to a pin weld on the upstream side of the magnet. Operators prepared MI-65 for an access.

The Plans for Monday

The plans for today are to allow the machines to conduct necessary maintenance. The Run Coordinator wants everyone to be finished by 3 PM, with beam to the neutrino experiments shortly there after. The TEV cooldown will begin on Tuesday, 1/4/11.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/05/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory